OpenStudy (anonymous):

\[\frac5z+\frac32_z= -1\]

OpenStudy (anonymous):

at the bottom there its 2z but i didnt know how to write it...

OpenStudy (kinggeorge):

First, multiply the \(5\over z\) by \(2\over2\). Now that you have a common denominator, add the two to get \[\frac{10+3}{2z}=-1\]

OpenStudy (kinggeorge):

Can you solve for \(z\) from there?

OpenStudy (anonymous):

Can I multiply both sides by 2z to get rid of the denominator 2z?

OpenStudy (kinggeorge):

That's exactly what you want to do.

OpenStudy (anonymous):

then divide both sides by -2 to isolate z, so z=-13/2..?

OpenStudy (kinggeorge):

Perfect.

OpenStudy (anonymous):

Woo! Thanks :D

OpenStudy (kinggeorge):

You're welcome.